%%% -*-BibTeX-*-
%%% Do NOT edit. File created by BibTeX with style
%%% ACM-Reference-Format-Journals [18-Jan-2012].

\begin{thebibliography}{00}

%%% ====================================================================
%%% NOTE TO THE USER: you can override these defaults by providing
%%% customized versions of any of these macros before the \bibliography
%%% command.  Each of them MUST provide its own final punctuation,
%%% except for \shownote{}, \showDOI{}, and \showURL{}.  The latter two
%%% do not use final punctuation, in order to avoid confusing it with
%%% the Web address.
%%%
%%% To suppress output of a particular field, define its macro to expand
%%% to an empty string, or better, \unskip, like this:
%%%
%%% \newcommand{\showDOI}[1]{\unskip}   % LaTeX syntax
%%%
%%% \def \showDOI #1{\unskip}           % plain TeX syntax
%%%
%%% ====================================================================

\ifx \showCODEN    \undefined \def \showCODEN     #1{\unskip}     \fi
\ifx \showDOI      \undefined \def \showDOI       #1{{\tt DOI:}\penalty0{#1}\ }
  \fi
\ifx \showISBNx    \undefined \def \showISBNx     #1{\unskip}     \fi
\ifx \showISBNxiii \undefined \def \showISBNxiii  #1{\unskip}     \fi
\ifx \showISSN     \undefined \def \showISSN      #1{\unskip}     \fi
\ifx \showLCCN     \undefined \def \showLCCN      #1{\unskip}     \fi
\ifx \shownote     \undefined \def \shownote      #1{#1}          \fi
\ifx \showarticletitle \undefined \def \showarticletitle #1{#1}   \fi
\ifx \showURL      \undefined \def \showURL       #1{#1}          \fi

\bibitem{iso23049}
{ISO/TR 23049:2018}. 2018.
\newblock {\em Road Vehicles: Ergonomic Aspects of External Visual
  Communication from Automated Vehicles to Other Road Users.}
\newblock Standard. International Organization for Standardization.
\newblock


\bibitem{ackermann2019experimental}
{Claudia Ackermann}, {Matthias Beggiato}, {Sarah Schubert}, {and} {Josef~F
  Krems}. 2019.
\newblock \showarticletitle{An experimental study to investigate design and
  assessment criteria: what is important for communication between pedestrians
  and automated vehicles?}
\newblock {\em Applied ergonomics\/}  {75} (2019), 272--282.
\newblock


\bibitem{aria2016investigation}
{Erfan Aria}. 2016.
\newblock Investigation of automated vehicle effects on driver s behavior and
  traffic performance.
\newblock   (2016).
\newblock


\bibitem{f015}
{Mercedes Benz}. 2015.
\newblock {Der Mercedes-Benz F 015 Luxury in Motion.}
\newblock
  \url{https://www.mercedes-benz.com/de/innovation/forschungsfahrzeug-f-015-luxury-in-motion/}.
    (2015).
\newblock
\newblock
\shownote{[Online; accessed: 07-DECEMBER-2019].}


\bibitem{colley2020towards}
{Mark Colley}, {Marcel Walch}, {Jan Gugenheimer}, {Ali Askari}, {and} {Enrico
  Rukzio}. 2020.
\newblock \showarticletitle{Towards Inclusive External Communication of
  Autonomous Vehicles for Pedestrians with Vision Impairments}. In {\em
  Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems}
  {\em (CHI ’20)}. Association for Computing Machinery, New York, NY, USA,
  1–14.
\newblock
\showISBNx{9781450367080}
\showDOI{%
\url{http://dx.doi.org/10.1145/3313831.3376472}}


\bibitem{colley2019including}
{Mark Colley}, {Marcel Walch}, {Jan Gugenheimer}, {and} {Enrico Rukzio}. 2019.
\newblock \showarticletitle{Including People with Impairments from the Start:
  External Communication of Autonomous Vehicles}. In {\em Proceedings of the
  11th International Conference on Automotive User Interfaces and Interactive
  Vehicular Applications: Adjunct Proceedings} {\em (AutomotiveUI ’19)}.
  Association for Computing Machinery, New York, NY, USA, 307–314.
\newblock
\showISBNx{9781450369206}
\showDOI{%
\url{http://dx.doi.org/10.1145/3349263.3351521}}


\bibitem{euco2}
{European Commission}.
\newblock {Road transport: Reducing CO2 emissions from vehicles}.
\newblock \url{https://ec.europa.eu/clima/policies/transport/vehicles_en}.
  (????).
\newblock
\newblock
\shownote{[Online; accessed: 31-JANUARY-2020].}


\bibitem{deb2019comparison}
{Shuchisnigdha Deb}, {Daniel~W. Carruth}, {Muztaba Fuad}, {Laura~M. Stanley},
  {and} {Darren Frey}. 2020.
\newblock \showarticletitle{Comparison of Child and Adult Pedestrian
  Perspectives of External Features on Autonomous Vehicles Using Virtual
  Reality Experiment}. In {\em Advances in Human Factors of Transportation},
  {Neville Stanton} (Ed.). Springer International Publishing, Cham, 145--156.
\newblock
\showISBNx{978-3-030-20503-4}


\bibitem{dey2018interface}
{Debargha Dey}, {Marieke Martens}, {Chao Wang}, {Felix Ros}, {and} {Jacques
  Terken}. 2018.
\newblock \showarticletitle{Interface Concepts for Intent Communication from
  Autonomous Vehicles to Vulnerable Road Users}. In {\em Adjunct Proceedings of
  the 10th International Conference on Automotive User Interfaces and
  Interactive Vehicular Applications} {\em (AutomotiveUI ’18)}. Association
  for Computing Machinery, New York, NY, USA, 82–86.
\newblock
\showISBNx{9781450359474}
\showDOI{%
\url{http://dx.doi.org/10.1145/3239092.3265946}}


\bibitem{fagnant2015preparing}
{Daniel~J Fagnant} {and} {Kara Kockelman}. 2015.
\newblock \showarticletitle{Preparing a nation for autonomous vehicles:
  opportunities, barriers and policy recommendations}.
\newblock {\em Transportation Research Part A: Policy and Practice\/}  {77}
  (2015), 167--181.
\newblock


\bibitem{fernandes2012platooning}
{Pedro Fernandes} {and} {Urbano Nunes}. 2012.
\newblock \showarticletitle{Platooning with IVC-enabled autonomous vehicles:
  Strategies to mitigate communication delays, improve safety and traffic
  flow}.
\newblock {\em IEEE Transactions on Intelligent Transportation Systems\/} {13},
  1 (2012), 91--106.
\newblock


\bibitem{RefWorks:doc:5cf7ad8de4b06bba938e0112}
{Evelyn Florentine}, {Mark~Adam Ang}, {Scott~Drew Pendleton}, {Hans Andersen},
  {and} {Marcelo~H. Ang}. 2016.
\newblock \showarticletitle{Pedestrian Notification Methods in Autonomous
  Vehicles for Multi-Class Mobility-on-Demand Service}. In {\em Proceedings of
  the Fourth International Conference on Human Agent Interaction} {\em (HAI
  ’16)}. Association for Computing Machinery, New York, NY, USA, 387–392.
\newblock
\showISBNx{9781450345088}
\showDOI{%
\url{http://dx.doi.org/10.1145/2974804.2974833}}


\bibitem{gonder2012analyzing}
{Jeffrey Gonder}, {Matthew Earleywine}, {and} {Witt Sparks}. 2012.
\newblock \showarticletitle{Analyzing vehicle fuel saving opportunities through
  intelligent driver feedback}.
\newblock {\em SAE International Journal of Passenger Cars-Electronic and
  Electrical Systems\/} {5}, 2012-01-0494 (2012), 450--461.
\newblock


\bibitem{hollander2019overtrust}
{Kai Holl\"{a}nder}, {Philipp Wintersberger}, {and} {Andreas Butz}. 2019.
\newblock \showarticletitle{Overtrust in External Cues of Automated Vehicles:
  An Experimental Investigation}. In {\em Proceedings of the 11th International
  Conference on Automotive User Interfaces and Interactive Vehicular
  Applications} {\em (AutomotiveUI ’19)}. Association for Computing
  Machinery, New York, NY, USA, 211–221.
\newblock
\showISBNx{9781450368841}
\showDOI{%
\url{http://dx.doi.org/10.1145/3342197.3344528}}


\bibitem{hudson2018pedestrian}
{Christopher~R. Hudson}, {Shuchisnigdha Deb}, {Daniel~W. Carruth}, {John
  McGinley}, {and} {Darren Frey}. 2019.
\newblock \showarticletitle{Pedestrian Perception of Autonomous Vehicles with
  External Interacting Features}. In {\em Advances in Human Factors and Systems
  Interaction}, {Isabel~L. Nunes} (Ed.). Springer International Publishing,
  Cham, 33--39.
\newblock
\showISBNx{978-3-319-94334-3}


\bibitem{inners2017beyond}
{Michael Inners} {and} {Andrew~L Kun}. 2017.
\newblock \showarticletitle{Beyond liability: Legal issues of human-machine
  interaction for automated vehicles}. In {\em Proceedings of the 9th
  International Conference on Automotive User Interfaces and Interactive
  Vehicular Applications}. ACM, Association for Computing Machinery, New York,
  NY, USA, 245--253.
\newblock


\bibitem{koo2015did}
{Jeamin Koo}, {Jungsuk Kwac}, {Wendy Ju}, {Martin Steinert}, {Larry Leifer},
  {and} {Clifford Nass}. 2015.
\newblock \showarticletitle{Why did my car just do that? Explaining
  semi-autonomous driving actions to improve driver understanding, trust, and
  performance}.
\newblock {\em International Journal on Interactive Design and Manufacturing
  (IJIDeM)\/} {9}, 4 (2015), 269--275.
\newblock


\bibitem{heise2019platooning}
{Stefan Krempl}. 2019.
\newblock {LKW-Platooning: Sicher, aber nicht so sparsam wie erwartet}.
\newblock
  \url{https://www.heise.de/newsticker/meldung/LKW-Platooning-Sicher-aber-nicht-so-sparsam-wie-erwartet-4419922.html}.
    (2019).
\newblock
\newblock
\shownote{[Online; accessed: 31-JANUARY-2020].}


\bibitem{liu2013opportunities}
{Chunhua Liu}, {KT Chau}, {Diyun Wu}, {and} {Shuang Gao}. 2013.
\newblock \showarticletitle{Opportunities and challenges of vehicle-to-home,
  vehicle-to-vehicle, and vehicle-to-grid technologies}.
\newblock {\it Proc. IEEE} {101}, 11 (2013), 2409--2427.
\newblock


\bibitem{locken2019should}
{Andreas L\"{o}cken}, {Carmen Golling}, {and} {Andreas Riener}. 2019.
\newblock \showarticletitle{How Should Automated Vehicles Interact with
  Pedestrians? A Comparative Analysis of Interaction Concepts in Virtual
  Reality}. In {\em Proceedings of the 11th International Conference on
  Automotive User Interfaces and Interactive Vehicular Applications} {\em
  (AutomotiveUI ’19)}. Association for Computing Machinery, New York, NY,
  USA, 262–274.
\newblock
\showISBNx{9781450368841}
\showDOI{%
\url{http://dx.doi.org/10.1145/3342197.3344544}}


\bibitem{RefWorks:doc:5cf8aa47e4b006bc06a90b0a}
{Victor~Malmsten Lundgren}, {Azra Habibovic}, {Jonas Andersson}, {Tobias
  Lagstr{\"o}m}, {Maria Nilsson}, {Anna Sirkka}, {Johan Fagerl{\"o}nn}, {Rikard
  Fredriksson}, {Claes Edgren}, {Stas Krupenia}, {and} {Dennis
  Salu{\"a}{\"a}r}. 2017.
\newblock \showarticletitle{Will There Be New Communication Needs When
  Introducing Automated Vehicles to the Urban Context?}. In {\em Advances in
  Human Aspects of Transportation}, {Neville~A. Stanton}, {Steven Landry},
  {Giuseppe Di~Bucchianico}, {and} {Andrea Vallicelli} (Eds.). Springer
  International Publishing, Cham, 485--497.
\newblock
\showISBNx{978-3-319-41682-3}


\bibitem{mahadevan2019av}
{Karthik Mahadevan}, {Elaheh Sanoubari}, {Sowmya Somanath}, {James~E. Young},
  {and} {Ehud Sharlin}. 2019.
\newblock \showarticletitle{AV-Pedestrian Interaction Design Using a Pedestrian
  Mixed Traffic Simulator}. In {\em Proceedings of the 2019 on Designing
  Interactive Systems Conference} {\em (DIS ’19)}. Association for Computing
  Machinery, New York, NY, USA, 475–486.
\newblock
\showISBNx{9781450358507}
\showDOI{%
\url{http://dx.doi.org/10.1145/3322276.3322328}}


\bibitem{mahadevan2018communicating}
{Karthik Mahadevan}, {Sowmya Somanath}, {and} {Ehud Sharlin}. 2018.
\newblock \showarticletitle{Communicating Awareness and Intent in Autonomous
  Vehicle-Pedestrian Interaction}. In {\em Proceedings of the 2018 CHI
  Conference on Human Factors in Computing Systems} {\em (CHI ’18)}.
  Association for Computing Machinery, New York, NY, USA, 1–12.
\newblock
\showISBNx{9781450356206}
\showDOI{%
\url{http://dx.doi.org/10.1145/3173574.3174003}}


\bibitem{matthiesen2018autonomous}
{Taggart Matthiesen}, {Jisi Guo}, {Sebastian Rolf~Johan Brannstrom}, {and}
  {Jess Garms}. 2018.
\newblock Autonomous vehicle notification system.
\newblock   (Dec.~11 2018).
\newblock
\newblock
\shownote{US Patent 10,152,892.}


\bibitem{millard2018pedestrians}
{Adam Millard-Ball}. 2018.
\newblock \showarticletitle{Pedestrians, autonomous vehicles, and cities}.
\newblock {\em Journal of planning education and research\/} {38}, 1 (2018),
  6--12.
\newblock


\bibitem{moore2019case}
{Dylan Moore}, {Rebecca Currano}, {G.~Ella Strack}, {and} {David Sirkin}. 2019.
\newblock \showarticletitle{The Case for Implicit External Human-Machine
  Interfaces for Autonomous Vehicles}. In {\em Proceedings of the 11th
  International Conference on Automotive User Interfaces and Interactive
  Vehicular Applications} {\em (AutomotiveUI ’19)}. Association for Computing
  Machinery, New York, NY, USA, 295–307.
\newblock
\showISBNx{9781450368841}
\showDOI{%
\url{http://dx.doi.org/10.1145/3342197.3345320}}


\bibitem{muller2016social}
{Lars M{\"u}ller}, {Malte Risto}, {and} {Colleen Emmenegger}. 2016.
\newblock \showarticletitle{The social behavior of autonomous vehicles}. In
  {\em Proceedings of the 2016 ACM International Joint Conference on Pervasive
  and Ubiquitous Computing: Adjunct}. ACM, New York, NY, USA, 686--689.
\newblock


\bibitem{RefWorks:doc:5cd92775e4b0487541989799}
{Andreas Sie\ss}, {Kathleen H\"ubel}, {Daniel Hepperle}, {Andreas Dronov},
  {Christian Hufnagel}, {Julia Aktun}, {and} {Matthias W\"olfel}. 2015.
\newblock \showarticletitle{Hybrid City Lighting-Improving Pedestrians' Safety
  through Proactive Street Lighting}. In {\em 2015 International Conference on
  Cyberworlds (CW)}. IEEE, New York, NY, USA, 46--49.
\newblock


\bibitem{stathopoulos2003induced}
{Fotis~G Stathopoulos} {and} {Robert~B Noland}. 2003.
\newblock \showarticletitle{Induced travel and emissions from traffic flow
  improvement projects}.
\newblock {\em Transportation Research Record\/} {1842}, 1 (2003), 57--63.
\newblock


\bibitem{walraven2016traffic}
{Erwin Walraven}, {Matthijs~TJ Spaan}, {and} {Bram Bakker}. 2016.
\newblock \showarticletitle{Traffic flow optimization: A reinforcement learning
  approach}.
\newblock {\em Engineering Applications of Artificial Intelligence\/}  {52}
  (2016), 203--212.
\newblock


\bibitem{wang2017hud}
{Chao Wang}, {Zhixiong Lu}, {Jacques Terken}, {and} {Jun Hu}. 2017.
\newblock \showarticletitle{HUD-AR: enhancing communication between drivers by
  affordable technology}. In {\em Proceedings of the 9th International
  Conference on Automotive User Interfaces and Interactive Vehicular
  Applications Adjunct}. ACM, ACM, New York, NY, USA, 249--250.
\newblock


\bibitem{RefWorks:doc:5ce536a5e4b076262b5839fa}
{Raphael Zimmermann} {and} {Reto Wettach}. 2017.
\newblock \showarticletitle{First step into visceral interaction with
  autonomous vehicles}. In {\em Proceedings of the 9th International Conference
  on Automotive User Interfaces and Interactive Vehicular Applications}. ACM,
  New York, NY, USA, 58--64.
\newblock


\end{thebibliography}

\documentclass{sigchi-ext}

\copyrightinfo{Presented at the workshop "Should I Stay or Should I Go? Automated Vehicles in the Age of Climate Change", April 25, 2020, Honolulu, HI, USA}

\usepackage[T1]{fontenc}
\usepackage{textcomp}
\usepackage[scaled=.92]{helvet} 
\usepackage{graphicx} 
\usepackage{balance}  
\usepackage{booktabs} 
\usepackage{ccicons}  
\usepackage{ragged2e} 
\usepackage{dirtytalk} 
\usepackage[utf8]{inputenc}  
\usepackage{todonotes} 
\usepackage{float} 
\usepackage{subcaption} 
\usepackage{txfonts}

\usepackage{url}
\makeatletter
\g@addto@macro{\UrlBreaks}{\UrlOrds}
\makeatother

\usepackage{multirow} 
\usepackage{tabularx} 

\usepackage[USenglish,UKenglish]{babel}
\usepackage[nodayofweek,level]{datetime}
\usepackage[USenglish]{isodate}

\usepackage{color} 

\usepackage{enumitem} 

\usepackage[official]{eurosym}
\usepackage{siunitx} 

\usepackage[normalem]{ulem}

\usepackage{microtype}        
\usepackage[all]{hypcap}      
\usepackage{ccicons}          

\usepackage{cite} 

\def\plaintitle{Towards Reducing Energy Waste through Usage of External Communication of Autonomous Vehicles} 
  \def\plainauthor{Mark Colley, Marcel, Walch, Enrico Rukzio}

\def\plainkeywords{Autonomous vehicles; self-driving vehicles; communication; external communication; fuel efficiency.}

\title{\plaintitle}

\numberofauthors{3}
\author{
  \alignauthor{%
    \textbf{Mark Colley}\\
    \affaddr{Institute of Media Informatics} \\
    \affaddr{Ulm University, Germany} \\
    \email{mark.colley@uni-ulm.de} } \vfil 
    \alignauthor{%
    \textbf{Marcel Walch}\\    
    \affaddr{Institute of Media Informatics} \\
    \affaddr{Ulm University, Germany} \\
    \email{marcel.walch@uni-ulm.de} }\vfil 
    \alignauthor{%
    \textbf{Enrico Rukzio}\\    
    \affaddr{Institute of Media Informatics} \\
    \affaddr{Ulm University, Germany} \\
    \email{enrico.rukzio@uni-ulm.de} }
    }

\definecolor{linkColor}{RGB}{6,125,233}
\hypersetup{%
  pdftitle={\plaintitle},
  pdfauthor={\plainauthor},
  pdfkeywords={\plainkeywords},
  bookmarksnumbered,
  pdfstartview={FitH},
  colorlinks,
  citecolor=black,
  filecolor=black,
  linkcolor=black,
  urlcolor=linkColor,
  breaklinks=true,
}

\begin{document}


\maketitle

\RaggedRight{} 

\begin{abstract}
Automated vehicles can implement strategies to drive with optimized fuel efficiency. Therefore, automated driving is seen as a major advancement in tackling climate change. 
However, with automated vehicles driving in cities and other areas rife with other road users such as human drivers, pedestrians, or cyclists, there is the potential for \say{stop-and-go} traffic. This would greatly diminish the possibility of automated vehicles to drive fuel-efficient.
We suggest using \textit{external communication of automated vehicles} to aid in ecological driving by providing clues to other road users to show the intent and therefore ultimately enable smoother traffic.

\end{abstract}

\keywords{\plainkeywords}

\begin{CCSXML}
<ccs2012>
<concept>
<concept_id>10003120.10003121.10003126</concept_id>
<concept_desc>Human-centered computing~HCI theory, concepts and models</concept_desc>
<concept_significance>300</concept_significance>
</concept>
</ccs2012>
\end{CCSXML}

\ccsdesc[500]{Human-centered computing~HCI theory, concepts and models}
\printccsdesc

\newpage
\section{Introduction}\label{sec:introduction}
Road traffic is a major contributor to the global emission of carbon dioxide. In Europe, for example, road transportation accounted for approximately 21\% of the carbon dioxide emissions~\cite{euco2}. Traffic congestion leads to increased usage of fuel~\cite{walraven2016traffic}. Therefore, ways to reduce congestion and, in general, to improve traffic flow, have to be found.
Autonomous traffic is expected to change traffic fundamentally~\cite{fagnant2015preparing, aria2016investigation}. 
Platooning is expected to improve traffic flow and fuel efficiency by reducing air resistance~\cite{fernandes2012platooning}. For this, real-time capable technologies are needed. First tests have shown the benefit of platooning while being not as efficient as hoped~\cite{heise2019platooning}. Smoothing of traffic flow in general was assessed as resulting in reduced emissions~\cite{stathopoulos2003induced}.

In recent years, the external communication of autonomous vehicles (AVs) has become a novel field of research. This research is targeted towards replacing interpersonal communication used today between human drivers and vulnerable road users (VRUs) towards improving traffic safety for people with disabilities~\cite{colley2020towards}. Various areas within this field have been explored: external communication modalities such as displays~\cite{RefWorks:doc:5cf7ad8de4b06bba938e0112}, LED strips~\cite{RefWorks:doc:5cf7ad8de4b06bba938e0112, RefWorks:doc:5cf8aa47e4b006bc06a90b0a}, movement patterns~\cite{RefWorks:doc:5ce536a5e4b076262b5839fa}, projections~\cite{ackermann2019experimental} auditory or tactile cues~\cite{mahadevan2018communicating}, as well as combinations thereof~\cite{mahadevan2018communicating} and enhancement of the infrastructure~\cite{RefWorks:doc:5cd92775e4b0487541989799}, trust challenges~\cite{hollander2019overtrust}, and (some) legal issues~\cite{inners2017beyond}.

While first tests with the mentioned technologies are promising, the hoped-for advances were not met. With the advancement of AVs, there is the fear of having even more miles travelled~\cite{fagnant2015preparing} on the street as these vehicles are able to look for parking spots in ever greater distance to the point of destination.

We introduce external communication of AVs towards other road users as one idea to smooth traffic flow. External communication of the intent~\cite{iso23049} can lead towards better understanding of the actions of AVs for pedestrians and therefore for fewer interruptions.
In this position paper, we outline classic traffic characteristics and how we believe external communication could influence these.

\section{External Communication of Autonomous Vehicles}
AVs do not need a human passenger to be present. Therefore, situations in which pedestrians and human drivers today rely on gestures or eye-contact could become very challenging. Recent research projects try to overcome these challenges via substitution of interpersonal communication with communication between AV and VRU (e.g.~\cite{locken2019should, colley2020towards, mahadevan2019av, deb2019comparison, dey2018interface}).
Müller et al.~\cite{muller2016social} already showed that AVs could be viewed as social actors in traffic. We report on related work in external communication of AVs to VRUs and other human drivers.

\subsection{External Communication Towards VRUs}
Various works have investigated communication between AVs and pedestrians. While there is still debate about the necessity of external communication~\cite{moore2019case}, many studies showed a benefit of external communication~\cite{colley2020towards, mahadevan2019av}. The current ISO standard~\cite{iso23049} advices to communicate intent towards VRUs instead of giving advice (\say{Safe to cross}) or even commands (\say{Cross}). However,other studies found that giving an unambiguous command was rated best~\cite{hudson2018pedestrian}.
The presented communication concepts have been categorized in various ways: by their complexity~\cite{locken2019should}, by their modality~\cite{mahadevan2018communicating, colley2020towards}, their information content~\cite{colley2020towards}, the locus of the communication~\cite{mahadevan2018communicating}, and their inclusiveness of people with visual impairments~\cite{colley2019including}. Different addressees have been compared, for example adults and children~\cite{deb2019comparison}.
Numerous technologies were proposed as described in the~\nameref{sec:introduction}.
However, this research only focused on a pedestrian who wants to cross a street.
\break
There is little literature about communication of AVs towards cyclists, however, there is a patent from Lyft~\cite{matthiesen2018autonomous} showing the AV communicating to the cyclist that it is \say{safe to cross}.

\subsection{External Communication towards Human Drivers}
Communicating the intent and explaining why an AV behaved the way it did was already investigated~\cite{koo2015did}. However, to our best knowledge, there is no current research on how an AV could communicate its intent to other drivers other than concepts from industry such as the F 015~\cite{f015} which, for example, indicates towards other human drivers to stop.
A video showcase presented by Wang et al.~\cite{wang2017hud} used a Head-Up Display (HUD) to show information about the other drivers and their intentions, e.g., driving to an airport. However, this was not in an autonomous context and other vehicles had to be equipped with visual markers. 
The aforementioned patent of Lyft~\cite{matthiesen2018autonomous} also showed that by using the side of the vehicle as a display, it can communicate that it is \say{yielding} at an intersection.

\section{Fuel Efficiency via External Communication}
Improving driver efficiency could lead (without AVs) to a reduction of about 20\%~\cite{gonder2012analyzing}. This would be achieved by avoiding stop-and-go, unnecessary idling and optimized acceleration and speed~\cite{gonder2012analyzing}.\break
With AVs, the same negative behavior has to be avoided. Optimized acceleration and speed will most likely be realized via algorithms, but stop-and-go and idling could be affected by humans.
Fuel efficiency could be decreased because of frequent slowing down and stops of AVs. In cities, a major reason for this could be people crossing in front of the vehicle. Millard-Ball already discussed this anticipated phenomenon as \say{crosswalk chicken}~\cite{millard2018pedestrians}. He claims that people will frequently cross streets in front of an AV as the vehicles will most likely be programmed to drive extra cautiously~\cite{millard2018pedestrians}.
This could be avoided through displays either by communicating early that a pedestrian can cross (cf.~\cite{locken2019should}) to avoid the need for a full stop or to to prevent pedestrians to cross too often (i.e., by displaying a warning). Millard-Ball~\cite{millard2018pedestrians} already states that an AV has to have some uncertainty whether it will stop to avoid too frequent crossings of pedestrians. Besides crossings, this could also be useful in crowded spaces such as parking lots or pedestrian areas.\break
With Avs, idling will become less. However, situations may arise in which a vehicle is waiting for a passenger (e.g., talking to other person). Depending on the settings, this could lead to unused energy in the form of heating. The vehicle could remind the passenger of this in a suitable way.

With regard to communication towards other human drivers, the intention of the AV can be communicated for example via external displays or via Vehicle-To-X (V2X) technology~\cite{liu2013opportunities}. A message showing that a (1) following vehicle has to slow down or that (2) a vehicle is about to turn (clearer than flashing lights) or (3) that a vehicle is yielding~\cite{matthiesen2018autonomous} can aid the other (manual) driver to avoid fuel inefficient maneuvers. To increase awareness towards fuel saving, an AV could also show on its display that it is currently driving in an eco-friendly way, therefore trying to persuade other to do the same, for example on a highway.

\section{Conclusion}
AVs are expected to reduce energy waste by driving more efficient~\cite{gonder2012analyzing}. However, there are unforeseeable events that could diminish this effect, such as pedestrians crossing frequently in front of these vehicles, causing a lot of stops~\cite{millard2018pedestrians}. We suggest to use external communication of AVs to reduce this. Additionally, we hope to encourage other human drivers to drive more efficiently by providing information about an eco-friendly driving style.
Moreover, we hope to spark a discussion about novel ways to reduce energy waste besides typical (technical or regulatory) approaches. Furthermore, we hope to gain knowledge from other domains and potentially start fruitful collaborations.


\section{Acknowledgements}
This work was conducted within the project 'Interaction between automated vehicles and vulnerable road users' (Intuitiver) funded by the Ministry of Science, Research and the Arts of the State of Baden-Württemberg.

\balance{} 

\bibliographystyle{SIGCHI-Reference-Format}
\bibliography{sample.bib}

\end{document}